OpenStudy (anonymous):

Which of the following equations is written in standard form? a. -2x=13 b. 3x+5y=-15 c. 1/2x+1/3y=1/5 d. -3x+13y=25. please i need help i dont understand what it means by standard form and it is on one my online tests i am sweating bullets right now can someone please help me out as soon as possible please. FIRST CORRECT ANSWER WILL RECEIVE A MEDAL!

OpenStudy (amistre64):

which one would you guess it is, then we can go from there

OpenStudy (anonymous):

i have no clue i am not understanding what they mean by standard form so i don't even no where to start.

OpenStudy (amistre64):

an since this is not a free answering service, trying to bribe a 'correct' answer by offering a medal is rather cheesy ... just saying

OpenStudy (amistre64):

standard from is like x+y = 4

OpenStudy (amistre64):

the site seems to be lagging a bit

OpenStudy (anonymous):

yeah i know it is chessy but i am desperate there is more than one question like that i want to know how to do it.

OpenStudy (amistre64):

standard from hates negative first numbers, and only likes to use integers

OpenStudy (amistre64):

so from this information, which one seems the most appropriate format?

OpenStudy (anonymous):

so a and d would be out then i am starting to think maybe c.

OpenStudy (amistre64):

a. -2x=13 b. 3x+5y=-15 c. 1/2x+1/3y=1/5 d. -3x+13y=25 a and d are bad correct b. 3x+5y=-15 c. 1/2x+1/3y=1/5 integers are not fractions ... standard form hates fractions and uses only integers

OpenStudy (anonymous):

so no negatives or fractions in the beginning so this one would be b. thanks for telling me that rule i think i can figure them out for myself now thanks again.

OpenStudy (amistre64):

:) thats the sites purpose ...

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks have a nice day.

OpenStudy (anonymous):

One more question i eliminated two answers already but i not sure which one of these would be in standard form 2x+3y=12 or 5x+5y10.